# EventMamba: Enhancing Spatio-Temporal Locality with State Space Models for Event-Based Video Reconstruction

### v2

@@ Introduction
Over the past decades, deep learning has achieved remarkable progress in computer vision (Ge, Fu, and Zha 2022; Ge et al. 2024; Zhang, Yang, and Wang 2023; Zhang et al. 2024; Zhang, Yang, and Hu 2023; Li et al. 2023; Shi et al. 2022; Peng et al. 2024; Wang et al. 2024), especially in event-based video reconstruction (EBVR) (Zhu et al. 2022; Rebecq et al. 2019; Cadena et al. 2023; Scheerlinck et al. 2020; Gallego et al. 2020). Despite these strides, EBVR algorithms still have room for improvement. Current methods typically utilize Convolutional Neural Networks (CNNs) or Transformers to reconstruct frames from event data. While CNNs focus on local details, often at the expense of global context, this can lead to increased susceptibility to noise and blurring, resulting in unclear visual outputs (Jang, McCormack, and Tong 2021). Conversely, Transformers excel in capturing extensive non-local information through their self-attention mechanisms (Weng, Zhang, and Xiong 2021; Xu et al. 2024). However, this approach scales quadratically with the input size, noted as $O(n^{2})$. This scaling is problematic for high-resolution data, such as the 1280$\times$720 output from Prophesee EVK4 cameras, which demands significant computational resources and challenges deployment on resource-limited devices.
@@
State Space Models (SSMs), first introduced in the S4 model (Gu, Goel, and Ré 2021), model global information more efficiently than CNNs or Transformers. S5 (Smith, Warrington, and Linderman 2022) reduced the complexity to linear levels using MIMO SSMs and parallel scanning. H3 (Mehta et al. 2022) added gating units to enhance expressiveness, enabling it to compete with Transformers in language modeling. Mamba (Gu and Dao 2023) introduced an input-adaptive mechanism, outperforming similarly-scaled Transformers in inference speed, throughput, and overall performance.
@@
In the training process, the symbol $\times$ denotes the Cartesian product. The parameters $(\xi_{l}^{h},\xi_{l}^{w})$ are considered as independently and identically distributed random variables sampled from the uniform distribution $\mathbb{U}(\mathfrak{R}_{s})$. Assuming the total number of Mamba layers is $N$, the random displacements $\{(\xi_{l}^{h},\xi_{l}^{w})\}_{l=0}^{N-1}$ are intentionally designed to be independent to ensure the maintenance of faithful locality and translation invariance at the layer level. Through this approach, although each individual layer considers all possible displacements, only one set of sampled displacements $\{(\xi_{l}^{h},\xi_{l}^{w})\}_{l=0}^{N-1}$ is required for each forward propagation, thus the training time remains consistent with that of fixed window partitioning. Following the RWO strategy, the event features are input into the VMB module and reshape to its original shape $E_{v}$.

### v3

@@ Introduction
Over the past decades, deep learning has achieved remarkable progress in computer vision (Ge, Fu, and Zha 2022; Ge et al. 2024; Zhang, Yang, and Wang 2023; Zhang et al. 2024; Zhang, Yang, and Hu 2023; Xiao et al. 2022; Li et al. 2023; Shi et al. 2022; Peng et al. 2024; Wang et al. 2024), especially in event-based video reconstruction (EBVR) (Zhu et al. 2022; Rebecq et al. 2019; Cadena et al. 2023; Scheerlinck et al. 2020; Gallego et al. 2020). Current methods typically utilize Convolutional Neural Networks (CNNs) or Transformers to reconstruct frames from event data. While CNNs focus on local details, often at the expense of global context, this can lead to increased susceptibility to noise and blurring, resulting in unclear visual outputs (Jang, McCormack, and Tong 2021). Conversely, Transformers excel in capturing extensive non-local information through their self-attention mechanisms (Weng, Zhang, and Xiong 2021; Xu et al. 2024). However, this approach scales quadratically with the input size, noted as $O(n^{2})$. This scaling issue is especially challenging for high-resolution data, like the 1280$\times$720 output from Prophesee EVK4 cameras, due to its high computational demands and the difficulty of deploying on devices with limited resources.
@@
State Space Models (SSMs), first introduced in the S4 model (Gu, Goel, and Ré 2021), model global information more efficiently than CNNs or Transformers. S5 (Smith, Warrington, and Linderman 2022) reduced the complexity to linear levels using MIMO SSMs and parallel scanning. H3 (Mehta et al. 2022) added gating units to enhance expressiveness, enabling it to compete with Transformers in language modeling. SSM-ViT (Zubić, Gehrig, and Scaramuzza 2024) significantly enhances performance and training speed at high frequencies by incorporating the SSMs into event-based vision. Mamba (Gu and Dao 2023) introduced an input-adaptive mechanism, outperforming similarly-scaled Transformers in inference speed, throughput, and overall performance.
@@
In the training process, the symbol $\times$ denotes the Cartesian product. The parameters $(\xi_{h}^{l},\xi_{w}^{l})$ are considered as independently and identically distributed random variables sampled from the uniform distribution $\mathbb{U}(\mathfrak{R}_{s})$. Assuming the total number of Mamba layers is $N$, the random displacements $\{(\xi_{h}^{l},\xi_{w}^{l})\}_{l=0}^{N-1}$ are intentionally designed to be independent to ensure the maintenance of faithful locality and translation invariance at the layer level. Through this approach, although each individual layer considers all possible displacements, only one set of sampled displacements $\{(\xi_{h}^{l},\xi_{w}^{l})\}_{l=0}^{N-1}$ is required for each forward propagation, thus the training time remains consistent with that of fixed window partitioning. Following the RWO strategy, the event features are input into the VMB module and reshape to its original shape $E_{v}$.